OpenStudy (anonymous):

If mass is equal to 89.87, the speed of light is 3x10 to the 8 and energy is e=mc squared, what is energy?

OpenStudy (wolf1728):

Here's a good calculator for that e=mc^2 stuff http://www.1728.org/einstein.htm but I'll also work it out for you.

OpenStudy (wolf1728):

First of all you have to get the units right. If we use mass of 89.87 grams and the speed of light is 29,979,245,800 centimeters per second (okay about 3 x 10^10 cm/sec) our answer will be in ergs When we use mass in kilograms, and speed of light equals 3 x 10^8 meters/second then the answer is in joules.

OpenStudy (wolf1728):

If you are using speed of light in meters per second, (3 x 10^8 m/sec) then we'll assume the 89.87 mass is in kilograms. E=mc^2 Energy (joules) = 89.87kg * (3 x 10^8 m/sec)^2 Energy = 89.87 * 9 x 10^16 Energy = 8.09 x 10^18 joules